OpenStudy (anonymous):

Solve this system of equations using the inverse of a matrix: x + y = 1 2x − z = 5 y − 3z = 6 Hint: AX = B and X = BA-1 x = –2.45, y = 1.73, and z = –1.57 x = 1.73, y = –2.45, and z = –1.57 x = 1.45, y = –2.17, and z = –2.15 x = –2.17, y = 1.45, and z = –2.15

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

oull the coeffs ad row reduce

OpenStudy (amistre64):

**pull the coeffs and row reduce

OpenStudy (anonymous):

Like 1x and 2x and 1y?

OpenStudy (amistre64):

x + y = 1 2x − z = 5 y − 3z = 6 fill in for all variables 1x + 1y + 0z= 1 2x + 0y − 1z = 5 0x + 1y − 3z = 6 now get rid of the variables or keep them, doesnt matter thay are immaterial row reduce using valid row operations

OpenStudy (anonymous):

(1.73)+(-2.45)= doesnt equal 1

OpenStudy (amistre64):

what are valid row operations?

OpenStudy (amistre64):

scalar multiplication, adding a scaled version of one row to the another .. seems like theres one more basic one that im forgetting

OpenStudy (anonymous):

row addition

OpenStudy (amistre64):

if you know how to work elimination process, its the same thing ...

OpenStudy (anonymous):

@amistre64 I think you're thinking of row interchanging?

OpenStudy (amistre64):

1x + 1y + 0z= 1 2x + 0y − 1z = 5 0x + 1y − 3z = 6 1 1 0 1 2 0 -1 5 0 1 -3 6 when we reduce this to the identity matrix on the left, the right column reduces to the solution (a,b,c) 1 0 0 a 0 1 0 b 0 0 1 c

OpenStudy (amistre64):

ah yes, swapping rows lol

OpenStudy (amistre64):

1 1 0 1 2 0 -1 5 0 1 -3 6 we want a 1 starting row1 and a 0 starting row2, and row3 ... we have a 1 already, and a 0 down below lets add a multiple of row1 to row2 such that thats 2 goes to 0; any ideas?

OpenStudy (amistre64):

its either this or try to determine an inverse .... this tends to be simpler in my book

OpenStudy (anonymous):

so you want a 1 starting row

OpenStudy (anonymous):

and a 0 on the second row

OpenStudy (amistre64):

for some value k, 1k + 2 = 0 is the thought process we want to occur

OpenStudy (anonymous):

wtf

OpenStudy (anonymous):

Im thinking about b

OpenStudy (amistre64):

if k=-2 we are set for it so -2(1 1 0 1) 2 0 -1 5 0 1 -3 6 -2 -2 0 -2 now add first 2 rows to get a new row 2 2 0 -1 5 0 1 -3 6 1 1 0 1 0 -2 -1 3 ; divide by -2 to get a 'leading' 1 0 1 -3 6 1 1 0 1 0 1 -.5 -1.5 0 1 -3 6 col2, wants to be 010 we have 111 add multiples of row2 to row1 and row3 .... that gets us there

OpenStudy (anonymous):

ok so multiply row 1 2 3

OpenStudy (anonymous):

row 1 is the second one you stated isn't it?

OpenStudy (amistre64):

row 1 is the first row. ive already worked it to the reduced form of: 1 1 0 1 0 1 -.5 -1.5 0 1 -3 6 -1 times row2 allows us to eliminate the [1] above it and the [1] below it

OpenStudy (amistre64):

1 1 0 1 -1( 0 1 -.5 -1.5) 0 1 -3 6 1 1 0 1 0 -1 .5 1.5 0 1 -3 6 now we can add row2 to row1 and row3 to get us the required 0 entries 1 0 .5 2.5 0 1 -.5 -1.5 0 0 -2.5 7.5 now we work the third row into submission, doing this by hand is of course for the birds lol

OpenStudy (anonymous):

mk lol

OpenStudy (amistre64):

divide the thrid row by -2.5 to get a leading 1; you see how we are building the left side into: 100 010 001 ???

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

1 0 .5 2.5 0 1 -.5 -1.5 0 0 1 -3 and of course what bites is we get all the way to the end and since none of your options has a -3 for z, it means i cant add or something ..... lets just avoid the errors and do this :) rref{{1 , 1 , 0 , 1},{2 , 0 , -1 , 5},{0 , 1 , -3 , 6}} http://www.wolframalpha.com/input/?i=rref%7B%7B1+%2C++1%2C+++0%2C+++++1%7D%2C%7B2%2C+++0%2C++-1%2C++++5%7D%2C%7B0%2C+++1%2C++-3%2C++++6%7D%7D

OpenStudy (anonymous):

I didn't think wolfram did matrices

OpenStudy (amistre64):

the wolf does just about everything, it getting the syntax correct thats key

OpenStudy (amistre64):

ive got to go put my leg up ... the pain is excruciating .. good luck

OpenStudy (anonymous):

wait what do I do next

OpenStudy (amistre64):

read the last column from the rref; its your solutions

OpenStudy (anonymous):

I got z

OpenStudy (anonymous):

ut there are 2 zs

OpenStudy (anonymous):

so it must be c

OpenStudy (amistre64):

1 1 0 1 0 1 -.5 -1.5 <-- i see an error, -1/-2 isnt -1/2 0 1 -3 6 1 1 0 1 0 1 .5 -1.5 0 1 -3 6 1 0 -.5 2.5 0 1 .5 -1.5 0 0 -3.5 7.5 1 0 -.5 2.5 0 1 .5 -1.5 0 0 1 -2.14 <-- thats better, now we are on track 2.14 - 1.5 out y is going to be positive so that leaves us with one option that has to work

OpenStudy (amistre64):

2.14(.5) - 1.5 1.07 - 1.5 yeah, its going to be c ...